A dealer has some hard candy worth $2 a pound and some worth $3 a pound. he wishes to make a mixture of 80 pounds that he can sell for $2.20 a pound. how many pounds of each kind should he use?
Answer by justasking(11)

I love these problems.
Make a picture of buckets. Under one, put "x", and in it put $2.
The second one should have a "80-x" under it (you can't have more than 80 pounds of hard candy), and in it write $3.
The third should have an "80" underneath it, and a $2.20 inside the bucket. Put a + between the first two, and a = between the the second and third buckets.
Now you can make your equation.
x*2 + 3(80-x)=80*2.20
2x + 240 - 3x = 176
2x + 240 = 176 + 3x
240 = 176 + x
64 = x
You need 64 pounds of the $2 candy.
80 - 64 = 16
You need 16 pounds of the $3 candy